OpenStudy (anonymous):

Why would you simplify 24x^4-8x^3+3/(4x-1)^2 into 24x^4-8x^3+3/(1-4x)^2. I'm confused as to what to do. I've currently found that dividing the whole fraction by -1 has the same effect.

OpenStudy (stamp):

I showed this in your previous thread. \[(4x-1)^2=(1-4x)^2\]\[(4x-1)^2=(4x-1)(4x-1)\]\[(4x-1)=(-1)(1-4x)\]\[(4x-1)^2=(-1)(1-4x)(-1)(1-4x)\]\[(-1)^2(1-4x)^2\]\[(4x-1)^2=(1-4x)^2\]